# How to set AdamW’s weight decay as you scale model and dataset size

### v1

@@ 4 Results
For all experiments, we implemented the model using PyTorch and performed optimization using its AdamW implementation. For all tasks, we did not use weight decay on the normalization layers. We used a cosine learning rate schedule, which reduced the learning rate by a factor of 10. All experiments were conducted on an internal cluster of Nvidia 1080ti/2080ti GPUs. Additional computation was expended on preliminary experiments. The results presented are the averages from three distinct random seeds. Comprehensive details on the hyperparameter search range and model specifications can be found in Appendix E.
4.1 The natural range of values for the AdamW EMA timescale
We considered a ResNet-18 (He et al., 2016) and ViT (Dosovitskiy et al., 2021) trained on CIFAR-10 (Krizhevsky, 2009) under a batchsize of 100, as such we have $M=50,000/100=500$. We trained the models with initial learning rates, $\eta$ of $\{10^{-4},10^{-3},10^{-2}\}$ for 200 epochs. For each $\eta$, we sweep $\tau_{\text{iter}}=1/(\eta\lambda)$ from $10^{-7}$ to $10^{-3}$ by changing $\lambda$.
The results are presented in Fig. 1: Broadly, we found that the optimal $\tau_{\text{epoch}}$ was usually between 1 and 200, which is clearly within our guidelines. While the optimal timescale was higher for the ViT with a learning rate of $10^{-4}$, the relevance of this finding is unclear as this learning rate is clearly suboptimal.
Figure 1: Optimal EMA timescale $\tau_{\text{epoch}}$ lies in a natural range. We trained a ResNet-18 (A) and a ViT (B) on CIFAR-10 using AdamW with cosine learning rate decay schedule under various initial learning rates $\eta$ (different lines) and different $\tau_{\text{epoch}}$ (x-axis; note we set $\lambda=1/(\eta M\tau_{\text{epoch}})$). We plotted various performance metrics after 200 epochs. The optimal values in terms of test metrics are highlighted with red crosses: For both models, the optimal $\tau_{\text{epoch}}$ lies in a natural range between 1 and 200 (shown by the gray dashed lines), except for $\eta=10^{-4}$ for ViT, which shows clearly suboptimal performance.
Moreover, we looked at large-scale LLM pretraining runs (Table 1) from Llama 1 (Touvron et al., 2023a), Llama 2 (Touvron et al., 2023b) and Stable-LM (Tow et al., 2023). Large-scale LLM pretraining runs are particularly interesting in our context because they are usually trained for only 1 or a few epochs. In that context, the EMA viewpoint suggests a narrow range of optimal weight decays: that $\tau_{\text{epoch}}$ should be neither much smaller nor much larger than $1$. While the use of a schedule, with smaller timescales initially and longer timescales at the end does complicate things slightly, the practical values chosen accord with those suggested by the EMA viewpoint. In particular, the initial timescales are smaller than 1 epoch, to ensure that very early contributions to the EMA may be forgotten, while the final timescale is around the total number of epochs, to ensure averaging over all datapoints. To obtain these timescales, we used the fact that that $\lambda=0.1$ for Llama 1, Llama 2, Stable LM, and Llama 1 + 2 decay the learning rate by a factor of 10, while Stable-LM decays to 4% of initial learning rate (Touvron et al., 2023a, b) and Stable-LM (Tow et al., 2023).
Table 1: EMA timescales in large-scale LLM pretraining runs
Model initial lr tokens/batch tokens epochs init $\tau_{\text{epoch}}$ final $\tau_{\text{epoch}}$
Llama 1 6.7B + 13B $3\times 10^{-4}$ 4 M 1 T 1 0.133 1.33
Llama 1 33B + 65B $1.5\times 10^{-4}$ 4 M 1.4 T 1 0.190 1.90
Llama 2 6.7B + 13B $3\times 10^{-4}$ 4 M 2 T 1 0.067 0.67
Llama 2 34B + 70B $1.5\times 10^{-4}$ 4 M 2 T 1 0.133 1.33
Stable-LM 3B $3.2\times 10^{-4}$ 4 M 1 T 4 0.125 3.13
@@ 4.2 Transferring weight decay across dataset sizes
Next, we study the relationship between optimal $\tau_{\text{epoch}}$ and the dataset size. Our prediction is that as dataset size increases, the optimal $\tau_{\text{epoch}}$ should remain broadly fixed. If we increase the dataset size with minibatch size fixed, then the number of minibatches in the dataset/an epoch, $M$, increases. Rearranging Eq. (8), that implies that $\tau_{\text{iter}}$ increases with dataset size, $\tau_{\text{iter}}=M\tau_{\text{epoch}}$. And rearranging $\tau_{\text{iter}}=1/(\eta\lambda)$,
$\lambda$ $=\frac{1}{\eta\tau_{\text{iter}}}=\frac{1}{\eta M\tau_{\text{epoch}}},$ (9)
tells us that if the optimal $\tau_{\text{epoch}}$ changes a little as the dataset size increases, we would predict that the optimal $\lambda$ should decrease as the dataset size increases.
To confirm this hypothesis, we trained a ResNet-18 and a ViT on ImageNet (Fig. 2). We used the $32\times 32$ downscaled version of ImageNet provided by Chrabaszcz et al. (2017) to ensure that we were able to perform the large number of training runs required to assess performance for different hyperparameter settings and dataset sizes. We trained the models on different subsets of ImageNet, where we randomly drew $80,160,320$, and $640$ samples from each of the $1,000$ classes except for the largest subset of $1.28M$ samples in total, where we randomly drew a subset from the whole dataset. We used a fixed batch size of 100 for all runs. We used an initial learning rate of $10^{-3}$ and swept $\lambda$ between $10^{-3}$ to $10^{3}$ and then plotted the performance metrics after 50 epochs. In the top row of Fig. 2A and Fig. 2B, we plotted the performance metrics vs. $\lambda$. Note that the optimal $\lambda$ decreases dramatically as the dataset size increases. In the bottom row of Fig. 2A and Fig. 2B, we plotted the performance metrics vs. $\tau_{\text{epoch}}$. Critically, the optimal $\tau_{\text{epoch}}$ is far more stable than the optimal $\lambda$.
@@
Figure 2: The optimal $\tau_{\text{epoch}}$ transfers across dataset sizes. We trained ResNet-18 (A) and ViT (B) on subsets of downsampled ImageNet of various sizes (lines of different colors) under different weight decay (dots on the lines) under a fixed batch size of 100. An initial learning rate of $10^{-3}$ is used with cosine decay scheduling. The performance metrics after 50 epochs are plotted against the weight decay $\lambda$ and the corresponding timescale $\tau_{\text{epoch}}$ computed with the initial learning rate. The dashed lines show the optimal $\tau_{\text{epoch}}$ at a subset size of $320,000$: In both models, the optimal $\tau_{\text{epoch}}$ is fairly stable across dataset sizes whereas the optimal $\lambda$ decreases dramatically as dataset size grows.
@@ 4.3 Transferring weight decay across model sizes
Figure 4: AdamW breaks the learning rate scaling of µP. Following the experiment setting in Yang et al. (2022), we trained a ResNet-18 with varying width factor $s$ (lines of different colors) under various base learning rates $\eta_{\text{base}}$ (x-axis) on CIFAR-10 (A) and a $320,000$ samples subset of ImageNet 32x32 (B). We then plotted the metrics after 200 (for CIFAR-10) and 50 (for ImageNet) epochs against $\eta_{\text{base}}$. The top row scales the hyperparameters using the direct µP approach (Eq. 11; i.e. fixed $\lambda$), while the bottom row scales the hyperparameters to ensure $\tau_{\text{iter}}$ is fixed (Eq. 13; $\lambda$ increases with model size). In both datasets, the direct approach breaks the stability of optimal $\eta_{\text{base}}$ in terms of test metrics due to changing the timescale whereas our scaling allows for consistent $\eta_{\text{base}}$ across model sizes.
Of course, it is also critical to understand how to modify the AdamW hyperparameters as we increase the model size. The most obvious approach to this issue is µP (Yang et al., 2022), which predicts that the optimal learning rate should decrease as $1/\text{fan\_in}$, by considering the behavior of the first few learning steps relative to the random initial weights. However, the theory behind µP did not consider how weight decay affects the learned weights in the later phase of the optimization.
Formally, we take $\eta_{\text{base}}$ and $\lambda_{\text{base}}$ as the learning rate and weight decay tuned on a smaller base model, with fan-in of $\text{fan\_in}_{\text{base}}$. Now, we consider scaling the model width by a factor of $s$,
$s$ $=\frac{\text{fan\_in}}{\text{fan\_in}_{\text{base}}}$ (10)
@@
$\eta$ $=\eta_{\text{base}}/s$ $\lambda$ $=\lambda_{\text{base}}.$ (11)
However, if we look at the implied EMA timescale,
@@
we see that this approach scales the EMA timescale with the model size. However, one of our hypotheses is that the optimal timescale should not vary with the model size. As such we conjecture that the scaling in Eq. (12) would break the transferability of optimal learning rate across model sizes.
How do we resolve this issue? We have two desiderata. First, from µP, we should scale $\eta$ with $1/\text{fan\_in}$. Second, we have the timescale $\tau_{\text{iter}}=1/(\eta\lambda)$ fixed. It might at first seem difficult to reconcile these two desiderata. Indeed, it is impossible in the usual setting where $\lambda$ is fixed. However, it is possible to reconcile them if we allow the weight decay, $\lambda$, to strengthen for larger models.
$\eta$ $=\eta_{\text{base}}/s$ $\lambda$ $=s\;\lambda_{\text{base}}.$ (13)
Thus,
@@
So the EMA timescale remains fixed as the model size increases.
@@
We begin by testing our prediction that the optimal EMA timescale was constant across model sizes, while the optimal weight decay was not, we trained a ResNet-18 of varying widths using µP’s codebase on $320,000$ samples subset of downscaled ImageNet, where we randomly drew $320$ samples from each class. We scaled the models’ widths by factors of 2, giving $s\in\{0.5,1,2\}$. We fixed $\eta_{\text{base}}=10^{-3}$ and used $\eta=\eta_{\text{base}}/s$ to change the learning rate with network size. We swept $\lambda_{\text{base}}$ from $10^{-2}$ to $10^{1}$. We then considered two ways of modifying the weight decay as we scaled the network and plotted the metrics after 50 epochs against weight decay. Fig. 3 (top row) leaves $\lambda$ fixed as the network size increases (Eq. 11). We can see the optimal $\lambda_{\text{base}}$ varies dramatically across network sizes. In contrast, Fig. 3 (bottom row) increases the weight decay as the network size increases (Eq. 13), as that leaves $\tau_{\text{iter}}$ unchanged (Eq. 14). It is evident that the optimal $\lambda_{\text{base}}$ and hence $\tau_{\text{iter}}$ is far more stable in this setting. Notably, while e.g. the optimal $\lambda_{\text{base}}$ for test loss on the bottom row does appear different from the others, if anything, this would indicate that the relationship between $s$ and $\lambda$ is even stronger than expected. In particular, this point indicates that even as we scale $\lambda$ with $s$, it the optimal $\lambda_{\text{base}}$ may still be increasing with $s$. That would seem to indicate that the relationship between $s$ and $\lambda_{\text{base}}$ might be superlinear (e.g. $\lambda_{\text{base}}\propto s^{\alpha}$, where $\alpha>1$), though we would need more, larger-scale experiments to definitively establish any such super-linearity, which is out-of-scope for the present work.
@@
Importantly, scaling the weight decay correctly has important implications even for the original µP predictions about how the optimal learning rate transfers across model sizes. Specifically, we trained a ResNet-18 on CIFAR-10 and again, on the 320K downscaled ImageNet subset, with a fixed $\lambda_{\text{base}}=1.0$ and swept $\eta_{\text{base}}$ from $10^{-5}$ to $10^{-2}$. In the top row of Fig. 4A and B, we use the direct µP scaling (Eq. 11), with a fixed weight decay. Remarkably, we found that the optimal base learning rate varied dramatically across model sizes. This indicates that µP scaling of the optimal learning rate breaks down in AdamW. Indeed, the original µP paper (Yang et al., 2022) did not examine AdamW, and this finding is confirmed by recent work on hyperparameter transfer in large-scale transformers (Lingle, 2024).
We hypothesized that this breakdown was due to the timescale changing with model size (Eq. 12), and thus, that we could restore the usual µP scaling of the learning rate by increasing the weight decay with model size (Eq. 13). To confirm this finding, in the bottom row of Fig. 4A and B, we increase the weight decay with model size (Eq. 13). We found that the optimal base learning rate was now far more stable when varying model sizes, confirming our hypothesis. Similar behavior was also observed when training ViTs of different widths on the ImageNet subset, the results are presented in Appendix C.
Finally, the size of the weight decay has implications for the magnitude of the learned weights, and hence whether the outputs $Wx$ remain $\mathcal{O}\left(1\right)$ which was one of the key desideratum used to derive µP (Yang et al., 2022). In particular, in AdamW the magnitude of the learned parameters, $W_{ij}$ scales with $1/\lambda$ (or $W_{ij}\sim 1/\lambda$). In our framework, this scaling arises because the quantity that AdamW takes the EMA of (i.e. $q_{t}$ in Eq.6) scales with $1/\lambda$), but this also arises in Kosson et al. (2023). As we propose $\lambda\sim\text{fan\_in}$, this suggests that the learned weights (not the initial weights) scale as, $W_{ij}\sim 1/\lambda\sim 1/\text{fan\_in}$. Thankfully, this is precisely the scaling we need to ensure that,
$y_{i}$ $=\sum_{j=1}^{\text{fan\_in}}W_{ij}x_{j}$ (15)
does not blow up or shrink as the model size increases (specifically, $y_{i}\sim 1$). In particular, the learned weights, $W_{ij}$ are a sum of many updates, $\Delta W_{ij}$. Thus, following the same reasoning as that for Eq. (2), we have $y_{i}\sim W_{ij}\times\text{fan\_in}$. Thus, to ensure that $y_{i}\sim 1$, we need $W_{ij}\sim 1/\text{fan\_in}$, which is precisely what we get by following our proposed scaling.

### v2

@@ 4 Scaling weight decay across model and dataset sizes
For all experiments, we implemented the model using PyTorch and performed optimization using its AdamW implementation. For all tasks, we did not use weight decay on the normalization layers. Experiments were conducted on an internal cluster of Nvidia TitanX/1080ti/2080ti/L40s GPUs, and only one GPU card was used at a time for each run. The results for image classification tasks are the averages from three distinct random seeds, for LLM pre-training tasks we only use one seed due to resource constraints. Comprehensive details on the hyperparameter search range and model specifications can be found in Appendix F.
@@ 4.1 Transferring weight decay across dataset sizes
Figure 1: The optimal $\tau_{\text{epoch}}$ transfers across dataset sizes. We trained ResNet-18 (A) and ViT (B) on subsets of downsampled ImageNet of various sizes (lines of different colors) under different weight decay (dots on the lines) under a fixed batch size of 100. An initial learning rate of $10^{-3}$ is used with cosine decay scheduling. The performance metrics after 100 epochs are plotted against the weight decay $\lambda$ and the corresponding timescale $\tau_{\text{epoch}}$ computed with the initial learning rate. The dashed lines show the optimal $\tau_{\text{epoch}}$ at a subset size of $320,000$: In both models, the optimal $\tau_{\text{epoch}}$ is fairly stable across dataset sizes whereas the optimal $\lambda$ decreases dramatically as dataset size grows.
Figure 2: For LLM pre-training, optimal weight decay shifts with dataset sizes but $\tau_{\text{epoch}}$ transfers. We trained a 124M NanoGPT on subsets of OpenWebText with different sizes (line colors) for 4 epochs under a fixed batch size, an initial learning rate of $6\times 10^{-4}$ and various $\lambda$ (dots on lines, varied in powers of 2). As dataset size increases, weight decay that gives optimal validation loss (red crosses) decreases, whereas $\tau_{\text{epoch}}$ (Eq. 10) is stable across scales.
First, we study the relationship between the optimal $\lambda$ and the training dataset size. Remember that the timescale, $\tau_{\text{iter}}$, intuitively measures the number of previous weight updates that we average over to get the weights. To understand just what proportion of the whole dataset we are averaging over, we propose to work with the timescale in epochs,
$\tau_{\text{epoch}}=\tau_{\text{iter}}/M=1/\left(\eta\lambda N/B\right),$ (10)
where $N$ is the number of training samples, $B$ is the batch size, and $M=N/B$ denotes the number of iterations in an epoch. Therefore, $\tau_{\text{epoch}}$ tells us how many epochs of past updates AdamW’s EMA averages over, where one epoch passes over all the data in the dataset.
Our prediction is that as dataset size increases, the optimal $\tau_{\text{epoch}}$ should remain approximately fixed. If we increase the dataset size with minibatch size fixed, then the number of minibatches in the dataset/an epoch, $M$, increases. Rearranging Eq. (10), that implies that $\tau_{\text{iter}}$ increases with dataset size, $\tau_{\text{iter}}=M\tau_{\text{epoch}}$. And rearranging $\tau_{\text{iter}}=1/(\eta\lambda)$,
$\lambda$ $=1/\left(\eta\tau_{\text{iter}}\right)=1/\left(\eta M\tau_{\text{epoch}}\right),$ (11)
tells us that if the optimal $\tau_{\text{epoch}}$ changes a little as the dataset size increases, we would predict that the optimal $\lambda$ should decrease as the dataset size increases. We confirmed this hypothesis on a ResNet-18 and a ViT trained on ImageNet (Fig. 1), along with NanoGPT (124M parameters) trained on OpenWebText (Fig. 2).
For the ImageNet experiments, we used the $32\times 32$ downscaled version of ImageNet provided by (Chrabaszcz et al., 2017) to ensure that we were able to perform the large number of training runs required to assess performance for different hyperparameter settings and dataset sizes. We trained the models on different subsets of ImageNet, where we randomly drew $80,160,320$, and $640$ samples from each of the $1,000$ classes except for the largest subset of $1.28M$ samples in total, where we randomly drew a subset from the whole dataset. We used a fixed batch size of 100 for all runs, an initial learning rate of $10^{-3}$, swept $\lambda$ from $10^{-3}$ to $10^{3}$, and then plotted the performance metrics after 100 epochs. The top row of Fig. 1A1B, shows performance vs. $\lambda$. Note that the optimal $\lambda$ decreases dramatically as the dataset size increases. The bottom row of Fig. 1A1B, shows performance metrics vs. $\tau_{\text{epoch}}$. Critically, the optimal $\tau_{\text{epoch}}$ is far more stable than the optimal $\lambda$. Here, we used a cosine schedule which decays to 0.1 of the initial learning rate. In the Appendix, we see similar patterns for a constant schedule, (Fig. 12) and cosine decay to zero (Fig. 13).
Next, we consider training a 12-layer NanoGPT with 124M parameters on subsets of OpenWebText (Gokaslan et al., 2019) of various sizes. We trained the model using the first 1/4, 1/2, and the whole dataset for approximately 4 epochs under a fixed batch size of 122880 tokens with a cosine learning rate decay schedule from $6\times 10^{-4}$ to $6\times 10^{-5}$. We swiped $\lambda$ between $2^{-8}$ to $2$, and plotted the final validation loss against optimal $\lambda$ and $\tau_{\text{epoch}}$ (Fig. 2), where we again observe that as dataset size grows, optimal $\lambda$ decreases but $\tau_{\text{epoch}}$ is more stable.
@@ 4.2 Transferring weight decay across model sizes
Figure 4: For LLM pre-training, optimal weight decay increases with model size whereas the optimal timescales transfer. We trained 8-layer GPTs on OpenWebText, with various widths (line colors), under two $\eta_{\text{base}}$, with actual learning rates for each width scaled following µP with $s=\tfrac{1024}{\textrm{width}}$. We plot the final validation loss against various $\lambda_{\text{base}}$ (dots on lines, varied by powers of $2$). We align all lines to the rightmost point for better visibility, where the numbers in the bracket denote the actual optimal validation loss at each width. We considered training under two parameterizations, if we keep $\lambda$ decoupled from $s$ then the optimal $\lambda_{\text{base}}$ shows a clear shift with model size (top rows), if $\lambda$ increases with $s$, optimal $\lambda_{\text{base}}$ becomes much stable across widths.
Figure 5: AdamW breaks the learning rate scaling of µP. Following the experiment setting in (Yang et al., 2022), we trained a ResNet-18 with varying width factor $s$ (lines of different colors) under various base learning rates $\eta_{\text{base}}$ (x-axis) on CIFAR-10 (A) and a $320,000$ samples subset of ImageNet 32x32 (B). We then plotted the metrics after 200 (for CIFAR-10) and 50 (for ImageNet) epochs against $\eta_{\text{base}}$. The top row scales the hyperparameters using the direct µP approach (Eq. 13; i.e. fixed $\lambda$), while the bottom row scales the hyperparameters to ensure $\tau_{\text{iter}}$ is fixed (Eq. 15; $\lambda$ increases with model size). In both datasets, the direct approach breaks the stability of optimal $\eta_{\text{base}}$ in terms of test metrics due to changing the timescale whereas our scaling allows for consistent $\eta_{\text{base}}$ across model sizes.
[…]
@@
$\eta$ $=\eta_{\text{base}}/s\quad\quad\quad\lambda=\lambda_{\text{base}},$ (13)
@@
However, the implied EMA timescale, now changes with model size, while we hypothesize that the optimal timescale should not vary with the model size. Therefore, we conjecture that the scaling in Eq. (14) would break the transferability of optimal learning rate across model sizes if we run the optimization for long enough (see Figure 5A B, top rows).
[…]
$\eta$ $=\eta_{\text{base}}/s\quad\quad\quad\lambda=s\;\lambda_{\text{base}},$ (15)
@@
We tested our conjecture that the optimal EMA timescale is constant across model sizes, while the optimal weight decay is not, by training a ResNet-18 of varying widths using µP’s codebase on $320,000$ samples subset of downscaled ImageNet, where we randomly drew $320$ samples from each class. We scaled the models’ widths by factors of 2, giving $s\in\{0.5,1,2\}$. We fixed $\eta_{\text{base}}=10^{-3}$ and used $\eta=\eta_{\text{base}}/s$ to change the learning rate with network size. We swept $\lambda_{\text{base}}$ from $10^{-2}$ to $10^{1}$. We then considered two ways of modifying the weight decay as we scaled the network and plotted the metrics after 50 epochs against weight decay. Fig. 3 (top row) leaves $\lambda$ fixed as the network size increases (Eq. 13). We can see the optimal $\lambda_{\text{base}}$ varies dramatically across network sizes. In contrast, Fig. 3 (bottom row) increases the weight decay as the network size increases (Eq. 15), as that leaves $\tau_{\text{iter}}$ unchanged (Eq. 16). It is evident that the optimal $\lambda_{\text{base}}$ and hence $\tau_{\text{iter}}$ is far more stable in this setting. Notably, while e.g. the optimal $\lambda_{\text{base}}$ for test loss on the bottom row does appear different from the others if anything, this would indicate that the relationship between $s$ and $\lambda$ is even stronger than expected. In particular, this point indicates that even as we scale $\lambda$ with $s$, the optimal $\lambda_{\text{base}}$ may still be increasing with $s$. That would seem to indicate that the relationship between $s$ and $\lambda_{\text{base}}$ might be superlinear (e.g. $\lambda_{\text{base}}\propto s^{\alpha}$, where $\alpha>1$), though we would need more, larger-scale experiments to definitively establish any such super-linearity, which is out-of-scope for the present work.
We conducted similar experiments for LLM pre-training, we considered 8-layer GPTs, with hidden states of widths in $\{256,512,1024\}$, achieved by changing the number of attention heads and fixing each head’s dimension. We considered two base learning rates, $3\times 10^{-4}$ and $3\times 10^{-5}$, trained for 250K iterations, and swiped lambda in the range $2^{-6}$ to $1$. The results are shown in Fig. 4, broadly, if we fix $\lambda$ irrespective of $s$ (top row, Eq. 13), optimal $\lambda$ increases with model sizes. However, as we scale $\lambda\propto s$ (bottom row, Eq. 14), the optimal $\lambda_{\text{base}}$ becomes consistent across widths, i.e. confirming that optimal timescale transfers.
@@
Importantly, scaling the weight decay correctly has important implications even for the original µP predictions about how the optimal learning rate transfers across model sizes. Specifically, we trained a ResNet-18 on CIFAR-10 and on 320K images downscaled from ImageNet, with a fixed $\lambda_{\text{base}}=1.0$ and swept $\eta_{\text{base}}$ from $10^{-5}$ to $10^{-2}$. In the top row of Fig. 5A and B, we use the standard µP scaling (Eq. 13), with a fixed weight decay. Remarkably, we found that the optimal base learning rate varied dramatically across model sizes. This indicates that µP scaling of the optimal learning rate breaks down in AdamW. Indeed, the original µP paper (Yang et al., 2022) did not examine AdamW, and this finding is confirmed by recent work on hyperparameter transfer in large-scale transformers (Lingle, 2024; Blake et al., 2024). We hypothesized that this breakdown was due to the timescale changing with model size (Eq. 14), and thus we could restore the usual µP scaling of the learning rate by increasing the weight decay with model size (Eq. 15) to fix the timescale when sweeping $\eta_{\text{base}}$. To confirm this finding, in the bottom row of Fig. 5A and B, we increase the weight decay with model size (Eq. 15). We found that the optimal base learning rate was now far more stable when varying model sizes, confirming our hypothesis. We observed similar behavior when training ViTs of different widths on the ImageNet subset, the results are presented in Appendix D, and follow-up work, citing us, observed similar behavior in large-scale training of transformers (Blake et al., 2024).
Lastly, we additionally considered constant and cosine decay to zero learning rate schedule for experiments Fig. 3 and Fig. 5, the results are shown in Fig. 1516 and Fig. 1718, respectively in the appendix.
[…]